OpenStudy (anonymous):

n=4; 2i and 3i are zeros; f(-2)=104

OpenStudy (anonymous):

you want a fourth degree polynomial with these zeros?

OpenStudy (anonymous):

start with \[a(x+2i)(x-2i)(x+3i)(x-3i)\] and multiply out

OpenStudy (anonymous):

yes how do i that?

OpenStudy (anonymous):

not as bad as it seems since \[(x+2i)(x-2i)=x^2+4\] and similarly \[(x+3i)(x-3i)=x^2+9\]

OpenStudy (anonymous):

so your last job is to write \[f(x)=a(x^2+4)(x^2+9)\] multiply that business out, and then to find \(a\) replace x by -2, set the result equal to 104 and solve for \(a\)

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

yes so far so good, then what?

OpenStudy (anonymous):

that is it

OpenStudy (anonymous):

o haha i did not see the last part, give me a second

OpenStudy (anonymous):

\[f(x)=a(x^4+13 x^2+36)\] and then \[f(-2)=104=a\left((-2)^4+13(-2)^2+36\right)\] solve for 'a'

OpenStudy (anonymous):

does a =1?

OpenStudy (anonymous):

i don't know i didn't compute it. hold on

OpenStudy (anonymous):

yes, because you get 104 for the stuff in the parentheses, so \(a=1\)

OpenStudy (anonymous):

ok the answer i am looking for say f(x)=. so would i just put f(x)=1?